# Edge-Cloud Collaborative Computing on Distributed Intelligence and Model Optimization: A Survey

### v4

@@ I-D Research Methodology and Organization
Recent analysis reveals significant research momentum in ECCC, as illustrated in Fig. 2, where publication trends and citation patterns demonstrate the evolving landscape across four interrelated domains (The statistics data is collected as of 2025/03/05). While traditional edge computing maintains consistent publication volumes with steadily increasing citation impact, cloud computing exhibits characteristics of a mature research area with decreasing publication numbers yet sustained high citation rates. Most notably, edge-cloud computing shows clear upward trajectories in both publications and citations, signaling growing research interest, whereas the specialized subfield of edge-cloud collaboration displays exponential citation growth peaking in 2024, indicating its emerging importance. Our systematic review methodology employed comprehensive literature searches across prominent scholarly databases including IEEE Xplore, ACM Digital Library, ScienceDirect, and Web of Science, utilizing carefully selected keywords such as ”edge-cloud computing”, ”distributed intelligence”, ”model optimization”, ”AI”, and ”machine learning”. Following retrieval, selected papers pertinent to ECCC were carefully reviewed and categorized based on their primary contributions aligned with the scope of this work.
@@ II-A Edge-Cloud Computing Paradigm
Distributed computing architectures optimize performance by positioning computational and storage resources closer to data sources, implementing a three-tier hierarchy encompassing cloud infrastructure, edge nodes, and end devices [47]. The cloud tier provides extensive computational and storage resources for large-scale data processing and model training, while the edge tier, comprising edge servers and gateways, offers localized processing to reduce latency [48]. End devices, such as smartphones and IoT sensors, generate and consume data at the network edge. Effective resource management and orchestration are crucial for efficient resource allocation across these tiers. Distributing intelligence across the network, a key characteristic of this paradigm, can be represented by a function $D(R_{i},L_{j})$, quantifying the proportion of resource $R_{i}$ allocated to layer $L_{j}$ [49]. Edge-cloud systems aim to optimize resource allocation, considering factors like application requirements and network conditions, to maximize a utility function $U(P,C,L)$ encompassing performance $P$, cost $C$, and latency $L$, subject to resource constraints [50]. Consequently, processing data closer to its source reduces latency and bandwidth consumption. Additionally, distributing workloads enhances scalability and fault tolerance, improving the system’s ability to handle increasing data volumes and withstand node failures [51]. Edge federation models further enable collaborative resource sharing, optimizing service delivery across a federated network [52], especially for mission-critical applications [53].
@@ II-B Distributed Intelligence Framework
Distributed intelligence frameworks optimize performance by strategically coordinating computational resources across heterogeneous infrastructure, leveraging both proximity-based edge processing and scalable cloud capabilities to address unique constraints and opportunities inherent in collaborative environments. The distributed intelligence framework in ECCC adheres to three key principles specifically adapted for edge-cloud deployment: task parallelism, data parallelism, and model parallelism.
Task parallelism in edge-cloud contexts executes different parts of complex AI tasks concurrently on separate devices, considering edge device limitations and cloud processing capabilities [54], represented as $T=\{t_{1}^{edge},t_{2}^{cloud},...,t_{n}^{hybrid}\}$, where tasks are strategically assigned based on latency requirements, computational complexity, and data locality. Data parallelism trains identical models on different data subsets distributed across multiple edge devices while maintaining privacy [11], expressed as $D=\{d_{1}^{local},d_{2}^{local},...,d_{m}^{local}\}$, where data remains decentralized to address privacy concerns and network bandwidth limitations. Model parallelism deploys different parts of large models on separate devices within the edge-cloud continuum [55], formulated as $M=\{m_{1}^{edge},m_{2}^{fog},...,m_{p}^{cloud}\}$, where model partitioning considers device capabilities, network latency, and energy constraints.
@@
Edge intelligence focuses on real-time processing with sub-10ms latency requirements [56], making it suitable for time-critical applications like autonomous driving and industrial control. Cloud intelligence handles complex computations and large-scale training using virtually unlimited resources [57], enabling sophisticated model development and global knowledge aggregation. Model partitioning strategies must balance workloads by considering model complexity, data size, network topology, and resource constraints specific to edge-cloud deployments [58]. Resource allocation algorithms dynamically assign computational, storage, and communication resources based on real-time demands, device capabilities, and edge-cloud topology [59]. Performance optimization techniques like model compression enhance efficiency across the distributed architecture by reducing communication overhead and enabling deployment on resource-constrained devices [60]. Additionally, fault tolerance mechanisms ensure system reliability despite potential edge device failures, network partitions, or cloud service interruptions, requiring redundancy strategies and graceful degradation protocols.
@@ II-C Resource Types and Constraints
Figure 3: Concepts of architecture. The diagram provides a holistic view of the technical scope covered in this survey, where the architecture integrates centralized cloud resources, decentralized edge nodes, and diverse mobile/edge devices. AI-enabled technologies include model compression, adaptation, and resource-aware optimization, whose components jointly support distributed intelligence and efficient model execution across multiple application scenarios.
Fundamental system resources, including computation, communication bandwidth, and storage capacity, present inherent constraints that significantly impact distributed architecture design and performance optimization across heterogeneous infrastructure [48]. Specifically, computation resources, denoted as $C_{k}$ for node $k$, represent processing power; communication bandwidth, $B_{ij}$, limits data transfer rates between nodes $i$ and $j$, directly affecting latency; and storage capacity, $S_{k}$, dictates the data volume node $k$ can store, influencing data-intensive applications [61]. Managing these constrained and heterogeneous resources across diverse devices, under dynamic workloads and with limited edge resources, especially energy for battery-powered devices [49], presents significant challenges.
Dynamic resource allocation addresses these challenges by adapting to workload fluctuations, aiming to find an optimal mapping of tasks $\mathcal{T}$ to nodes $\mathcal{N}$ that minimizes a cost function (e.g., latency) subject to resource constraints [62]. Mathematically, this involves minimizing $\sum_{t\in\mathcal{T}}Cost(M(t),t)$ subject to $\forall n\in\mathcal{N}:\sum_{t:M(t)=n}R(t)\leq C_{n}$, where $Cost(M(t),t)$ is the cost of executing task $t$ on node $M(t)$, and $R(t)$ is the resource requirement of task $t$. In addition, energy-efficient offloading minimizes energy consumption, $E_{ij}(t)=E_{comm}(t,B_{ij})+E_{comp}(t,C_{j})$, by strategically offloading tasks, where $E_{comm}$ and $E_{comp}$ represent communication and computation energy, respectively [39]. Efficient edge storage management, leveraging techniques like data caching and prefetching [63], is also crucial, especially for data-intensive applications, alongside integrating QoS considerations [64].
@@ III Model Optimization for Edge-Cloud Systems
Collaborative computing environments present distinctive optimization challenges that fundamentally differ from traditional centralized or purely cloud-based approaches. Distributed environments impose specific constraints that significantly impact optimization strategies: (1) Extreme resource heterogeneity, edge devices range from smartphones with limited processing power to industrial gateways with substantial computational capabilities, necessitating adaptive optimization techniques that can dynamically scale across this spectrum; (2) Intermittent connectivity, network links between edge and cloud are often unreliable, requiring optimization methods that can function during disconnection periods and efficiently synchronize when connectivity is restored; (3) Real-time processing requirements, many edge applications demand ultra-low latency responses (<< 10ms), making traditional batch optimization approaches unsuitable; (4) Privacy-preserving constraints, sensitive data often cannot leave edge devices, requiring optimization techniques that work with limited data visibility; (5) Energy limitations, battery-powered edge devices necessitate energy-aware optimization that balances computational efficiency with power consumption. Consequently, model optimization techniques must be specifically adapted to address these edge-cloud specific constraints, moving beyond general-purpose optimization to develop specialized approaches that leverage the unique characteristics of the edge-cloud continuum while mitigating its inherent limitations.
@@ III-A Model Compression
Resource-constrained environments demand sophisticated model compression techniques to enable deployment of complex AI models while maintaining acceptable performance levels despite limited computational power, memory, and energy resources. Advanced compression approaches [2, 65, 66] have emerged as critical solutions to reduce model size and computational demands while preserving performance. Table II provides a comparison of recent approaches across three primary compression categories: pruning, quantization, and knowledge distillation.
Model compression techniques in ECCC systems can be systematically categorized based on their compression mechanisms and deployment characteristics: (1) Structural modification approaches (pruning) that reduce model complexity by eliminating redundant parameters or architectural components, particularly effective for over-parameterized networks commonly deployed in edge scenarios; (2) Numerical precision optimization (quantization) that maintains model structure while reducing parameter precision requirements, enabling efficient deployment on specialized edge hardware with quantization support; (3) Knowledge transfer and distillation that creates compact models through supervised learning from larger teacher models, providing architectural flexibility and optimal accuracy-efficiency trade-offs for edge deployment. The classification framework reflects both the underlying technical mechanisms and practical deployment considerations essential for edge-cloud environments.
Each compression category exhibits distinct advantages and limitations for edge-cloud deployment scenarios. Pruning techniques excel in reducing model size and computational complexity by eliminating redundant parameters, making them particularly effective for convolutional neural networks where structured pruning can remove entire channels or layers. However, pruning may require fine-tuning to recover accuracy losses and can be sensitive to the pruning ratio selection [67]. Quantization methods provide significant memory reduction and accelerated inference by reducing parameter precision, with minimal accuracy degradation when carefully implemented. They are especially suitable for deployment on specialized hardware with quantization support, though they may face challenges with extremely low-bit quantization or complex model architectures [68]. Knowledge distillation offers the unique advantage of training compact models from scratch while leveraging teacher model knowledge, enabling flexible student model architectures. However, it requires access to both teacher models and training data, making it computationally intensive during the training phase [69].
@@
Notably, pruning techniques demonstrate significant improvements in inference latency and energy efficiency by strategically eliminating less important parameters [70], as further detailed in Sec. III-A1. Meanwhile, quantization methods, explored in Sec. III-A2, reduce parameter precision requirements, with implementations achieving remarkable energy efficiency in healthcare applications. Knowledge distillation approaches, discussed in Sec. III-A3, enable smaller models to effectively mimic larger counterparts, with recent innovations specifically optimized for efficient LLM inference across edge-cloud boundaries. Across these diverse techniques, a common pattern emerges wherein most successful implementations leverage both edge and cloud resources collaboratively, highlighting the synergistic nature of modern compression strategies in distributed computing environments.
TABLE II: Summary of model compression techniques.
@@
TSCF [67] 2024 IEEE IoTJ P Edge-Cloud Systems Inference Latency
Hybrid SD [71] 2024 ArXiv P Image Synthesis Parameter Efficiency, Image Quality
PSFed [72] 2024 IEEE JSAC P Satellite Edge Cloud Latency, Energy, Privacy
3CE2C [73] 2024 ASC Q Healthcare/ECG Accuracy, Energy Efficiency
CDC [74] 2020 IJCAI P Edge-Cloud Systems Bandwidth Efficiency
QuAsyncFL [68] 2024 IEEE IoTJ Q AIoT Communication Efficiency
Zhang et al. [75] 2025 IEEE JSAC Q Aerial Edge Computing Mean Average Precision (mAP)
DTS-KD [65] 2024 IEEE IoTJ K Industrial IoT F1 Score, Accuracy, Precision, Recall
Wang et al. [69] 2024 CVPR K Multimodal LLMs Generalization Performance
Wu et al. [76] 2024 IEEE TII P Aviation Manufacturing Edge Detection Accuracy
FedAgg [77] 2024 INFOCOM K End-Edge-Cloud Computing Model Accuracy, Convergence Rate
Yang et al. [6] 2025 ADHOC K IoT Traffic Identification Classification Accuracy, Resource Efficiency
GKT [78] 2024 ACL K LLM Inference Inference Speed, Accuracy, Cost
OTO [79] 2024 IEEE TPDS K Reinforcement Learning Convergence Speed, Reward
Zhang et al. [80] 2024 ESWA K Manufacturing Processes Fault Detection Accuracy
@@ III-A1 Pruning
Parameter elimination strategies reduce model complexity by removing redundant components, enabling efficient AI deployment on resource-constrained devices through targeted removal of less critical weights or architectural elements [71]. We explore various strategies targeting different architectural aspects, such as pruning entire neurons or filters based on performance contribution, as in the two-phase split computing framework (TSCF) [67], or focusing on individual weights. Applied during or after training, dynamic and static pruning balance accuracy, and computational cost. For instance, Hybrid SD [71] employs structural pruning of the Stable Diffusion Model’s U-Net to maintain image quality on edge devices, while the adaptive pruning-split federated learning (PSFed) method [72] optimizes semantic coder updates in a SemCom-assisted SEC framework, reducing training delay and energy consumption.

### v5

@@ I-D Research Methodology and Organization
Recent analysis reveals significant research momentum in ECCC, as illustrated in Fig. 2, where publication trends and citation patterns demonstrate the evolving landscape across four interrelated domains (The statistics data is collected as of 2025/03/05). While traditional edge computing maintains consistent publication volumes with steadily increasing citation impact, cloud computing exhibits characteristics of a mature research area with decreasing publication numbers yet sustained high citation rates. Most notably, edge-cloud computing shows clear upward trajectories in both publications and citations, signaling growing research interest, whereas the specialized subfield of edge-cloud collaboration displays exponential citation growth peaking in 2024, indicating its emerging importance. Our systematic review methodology employed comprehensive literature searches across prominent scholarly databases including IEEE Xplore, ACM Digital Library, ScienceDirect, and Web of Science, utilizing carefully selected keywords such as “edge-cloud computing”, “distributed intelligence”, “model optimization”, “AI”, and “machine learning”. Following retrieval, selected papers pertinent to ECCC were carefully reviewed and categorized based on their primary contributions aligned with the scope of this work.
@@ II-A Edge-Cloud Computing Paradigm
Distributed computing architectures optimize performance by positioning computational and storage resources closer to data sources, implementing a three-tier hierarchy encompassing cloud infrastructure, edge nodes, and end devices [vo2022edge]. Within this hierarchy, the cloud tier provides extensive computational and storage resources for large-scale data processing and model training, while the edge tier, comprising edge servers and gateways, offers localized processing to reduce latency [tocze2018taxonomy]. At the periphery, end devices such as smartphones and IoT sensors generate and consume data at the network edge, necessitating effective resource management and orchestration for efficient resource allocation across these tiers. Notably, distributing intelligence across the network, a key characteristic of this paradigm, can be represented by a function $D(R_{i},L_{j})$, quantifying the proportion of resource $R_{i}$ allocated to layer $L_{j}$ [hamm2019edge]. Edge-cloud systems aim to optimize resource allocation, considering factors like application requirements and network conditions, to maximize a utility function $U(P,C,L)$ encompassing performance $P$, cost $C$, and latency $L$, subject to resource constraints [tosic2019blockchainbased]. Consequently, processing data closer to its source reduces latency and bandwidth consumption. Additionally, distributing workloads enhances scalability and fault tolerance, improving the system’s ability to handle increasing data volumes and withstand node failures [cao2020edge]. Edge federation models further enable collaborative resource sharing, optimizing service delivery across a federated network [ramu2023edge], especially for mission-critical applications [skarin2018missioncritical].
@@ II-B Distributed Intelligence Framework
Distributed intelligence frameworks optimize performance by strategically coordinating computational resources across heterogeneous infrastructure, leveraging both proximity-based edge processing and scalable cloud capabilities to address unique constraints and opportunities inherent in collaborative environments. Within this framework, the distributed intelligence model in ECCC adheres to three key principles specifically adapted for edge-cloud deployment: task parallelism, data parallelism, and model parallelism.
Task parallelism in edge-cloud contexts executes different parts of complex AI tasks concurrently on separate devices, considering edge device limitations and cloud processing capabilities [wang2022ace], represented as $T=\{t_{1}^{edge},t_{2}^{cloud},...,t_{n}^{hybrid}\}$, where tasks are strategically assigned based on latency requirements, computational complexity, and data locality. Data parallelism trains identical models on different data subsets distributed across multiple edge devices while maintaining privacy [yao2023edgecloud], expressed as $D=\{d_{1}^{local},d_{2}^{local},...,d_{m}^{local}\}$, where data remains decentralized to address privacy concerns and network bandwidth limitations. Model parallelism deploys different parts of large models on separate devices within the edge-cloud continuum [kim2024dnn], formulated as $M=\{m_{1}^{edge},m_{2}^{fog},...,m_{p}^{cloud}\}$, where model partitioning considers device capabilities, network latency, and energy constraints.
@@
Edge intelligence focuses on real-time processing with sub-10ms latency requirements [asim2020review], making it suitable for time-critical applications like autonomous driving and industrial control. Cloud intelligence handles complex computations and large-scale training using virtually unlimited resources [mungoli2023scalable], enabling sophisticated model development and global knowledge aggregation. Model partitioning strategies must balance workloads by considering model complexity, data size, network topology, and resource constraints specific to edge-cloud deployments [banitalebi-dehkordi2021autosplit]. Resource allocation algorithms dynamically assign computational, storage, and communication resources based on real-time demands, device capabilities, and edge-cloud topology [alvar2021paretooptimal]. Performance optimization techniques like model compression enhance efficiency across the distributed architecture by reducing communication overhead and enabling deployment on resource-constrained devices [wang2024socialized]. Additionally, fault tolerance mechanisms ensure system reliability despite potential edge device failures, network partitions, or cloud service interruptions, requiring redundancy strategies and graceful degradation protocols.
@@ II-C Resource Types and Constraints
Figure 3: Concepts of architecture, wherein this diagram provides a holistic view of the technical scope covered in this survey, where the architecture integrates centralized cloud resources, decentralized edge nodes, and diverse mobile/edge devices while AI-enabled technologies including model compression, adaptation, and resource-aware optimization jointly support distributed intelligence and efficient model execution across multiple application scenarios.
Fundamental system resources, including computation, communication bandwidth, and storage capacity, present inherent constraints that significantly impact distributed architecture design and performance optimization across heterogeneous infrastructure [tocze2018taxonomy]. Specifically, computation resources, denoted as $C_{k}$ for node $k$, represent processing power; communication bandwidth, $B_{ij}$, limits data transfer rates between nodes $i$ and $j$, directly affecting latency; and storage capacity, $S_{k}$, dictates the data volume node $k$ can store, influencing data-intensive applications [luo2021resource]. Managing these constrained and heterogeneous resources across diverse devices, under dynamic workloads and with limited edge resources, especially energy for battery-powered devices [hamm2019edge].
Dynamic resource allocation addresses these challenges by adapting to workload fluctuations, aiming to find an optimal mapping of tasks $\mathcal{T}$ to nodes $\mathcal{N}$ that minimizes a cost function (e.g., latency) subject to resource constraints [rublein2024improved]. Mathematically, this involves minimizing $\sum_{t\in\mathcal{T}}Cost(M(t),t)$ subject to $\forall n\in\mathcal{N}:\sum_{t:M(t)=n}R(t)\leq C_{n}$, where $Cost(M(t),t)$ is the cost of executing task $t$ on node $M(t)$, and $R(t)$ is the resource requirement of task $t$. In addition, energy-efficient offloading minimizes energy consumption, $E_{ij}(t)=E_{comm}(t,B_{ij})+E_{comp}(t,C_{j})$, by strategically offloading tasks, where $E_{comm}$ and $E_{comp}$ represent communication and computation energy, respectively [wang2019edge]. Efficient edge storage management, leveraging techniques like data caching and prefetching [sonbol2020edgekv], is also crucial, especially for data-intensive applications, alongside integrating QoS considerations [farhoudi2023qosaware].
@@ III Model Optimization for Edge-Cloud Systems
Model Optimization (§III) Model Compression (§III-A) Model Adaptation (§III-B) Resource-Aware Opt. (§III-C) Neural Arch. Search (§III-D) Pruning Quantization Knowledge Distillation Transfer Learning Federated Learning Continual Learning Split Learning Dynamic Model Sel. Model Splitting RL-based NAS Evolutionary NAS Gradient-based NAS [ko2024twophase, yan2024hybrid, zheng2024semantic] [dong2020cdc, liu2024quasyncfl, chen2024implementing] [yao2024gkt, wu2024agglomerative, wang2024clouddevice] [li2024transfer, wang2024cloudedgea, sharma2024deep] [yu2024fedcae, le2024sfetec, zhao2024flexiblefl] [li2024digital, lin2024online, zhang2024crossfcl] [lin2024efficient, ahmed2024heterogeneous, kim2021splitbridge] [carpio2020engineering, rublein2024improved, wang2024efficient] [guo2024madrlom, hua2024energyefficient, wang2024cooperative] [goudarzi2023distributed, bensalem2023scaling, wen2023adaptivenet] [he2024efficient, chen2024fast, materwala2022performance] [yao2020efficient, dutta2023searchtime, li2021appealnet]
Figure 4: Taxonomy of model optimization.
To systematically address these challenges, this section presents a comprehensive taxonomy of model optimization techniques tailored for ECCC, as visually summarized in Fig. 4. Our classification is structured around four fundamental pillars that represent distinct yet complementary approaches to enhancing model efficiency and performance in distributed settings. Model compression (Sec. III-A) focuses on reducing model size and computational complexity. Model adaptation (Sec. III-B) enables models to adjust to new data or environments. Resource-aware model optimization (Sec. III-C) dynamically tailors models to hardware and network constraints. Finally, neural architecture search (Sec. III-D) automates the discovery of optimal model designs. Collectively, this taxonomy provides a structured framework for navigating the diverse landscape of optimization strategies.
@@ III-A Model Compression
Model compression techniques are essential for deploying complex AI models in resource-constrained environments by reducing their size and computational demands [tang2024anomaly, vuruma2024cloud]. In line with established literature, this survey categorizes these techniques into three widely recognized pillars: pruning, quantization, and knowledge distillation [wang2024endedgecloud]. Notably, this classification is justified by the distinct mechanism each approach employs. Pruning involves structural modification to eliminate redundant model parameters or components. Quantization focuses on numerical precision optimization by reducing the bit-width of parameters. Knowledge distillation facilitates knowledge transfer from a large “teacher” model to a smaller “student” model. As summarized in Table III, each category offers unique trade-offs between compression ratio, accuracy, and computational overhead, making them suitable for different ECCC scenarios.
Each compression category exhibits distinct advantages and limitations for edge-cloud deployment scenarios. Pruning techniques excel in reducing model size and computational complexity by eliminating redundant parameters, making them particularly effective for convolutional neural networks where structured pruning can remove entire channels or layers. However, pruning may require fine-tuning to recover accuracy losses and can be sensitive to the pruning ratio selection [ko2024twophase]. Quantization methods provide significant memory reduction and accelerated inference by reducing parameter precision, with minimal accuracy degradation when carefully implemented. They are especially suitable for deployment on specialized hardware with quantization support, though they may face challenges with extremely low-bit quantization or complex model architectures [liu2024quasyncfl]. Knowledge distillation offers the unique advantage of training compact models from scratch while leveraging teacher model knowledge, enabling flexible student model architectures. However, it requires access to both teacher models and training data, making it computationally intensive during the training phase [wang2024clouddevice].
@@
Notably, pruning techniques demonstrate significant improvements in inference latency and energy efficiency by strategically eliminating less important parameters [wu2022serving], as further detailed in Sec. III-A1. Meanwhile, quantization methods, explored in Sec. III-A2, reduce parameter precision requirements, with implementations achieving remarkable energy efficiency in healthcare applications. Knowledge distillation approaches, discussed in Sec. III-A3, enable smaller models to effectively mimic larger counterparts, with recent innovations specifically optimized for efficient LLM inference across edge-cloud boundaries. Across these diverse techniques, a common pattern emerges wherein most successful implementations leverage both edge and cloud resources collaboratively, highlighting the synergistic nature of modern compression strategies in distributed computing environments.
TABLE III: Summary of model compression techniques.
@@
TSCF [ko2024twophase] 2024 IEEE IoTJ P Edge-Cloud Systems Inference Latency
Hybrid SD [yan2024hybrid] 2024 ArXiv P Image Synthesis Parameter Efficiency, Image Quality
PSFed [zheng2024semantic] 2024 IEEE JSAC P Satellite Edge Cloud Latency, Energy, Privacy
3CE2C [chen2024implementing] 2024 ASC Q Healthcare/ECG Accuracy, Energy Efficiency
CDC [dong2020cdc] 2020 IJCAI P Edge-Cloud Systems Bandwidth Efficiency
QuAsyncFL [liu2024quasyncfl] 2024 IEEE IoTJ Q AIoT Communication Efficiency
Zhang et al. [zhang2025large] 2025 IEEE JSAC Q Aerial Edge Computing Mean Average Precision (mAP)
DTS-KD [tang2024anomaly] 2024 IEEE IoTJ K Industrial IoT F1 Score, Accuracy, Precision, Recall
Wang et al. [wang2024clouddevice] 2024 CVPR K Multimodal LLMs Generalization Performance
Wu et al. [wu2024systematic] 2024 IEEE TII P Aviation Manufacturing Edge Detection Accuracy
FedAgg [wu2024agglomerative] 2024 INFOCOM K End-Edge-Cloud Computing Model Accuracy, Convergence Rate
Yang et al. [yang2025growthadaptive] 2025 ADHOC K IoT Traffic Identification Classification Accuracy, Resource Efficiency
GKT [yao2024gkt] 2024 ACL K LLM Inference Inference Speed, Accuracy, Cost
OTO [zeng2024offlinetransferonline] 2024 IEEE TPDS K Reinforcement Learning Convergence Speed, Reward
Zhang et al. [zhang2024cloud] 2024 ESWA K Manufacturing Processes Fault Detection Accuracy
@@ III-A1 Pruning
Parameter elimination strategies reduce model complexity by removing redundant components, enabling efficient AI deployment on resource-constrained devices through targeted removal of less critical weights or architectural elements [yan2024hybrid]. We explore various strategies targeting different architectural aspects, such as pruning entire neurons or filters based on performance contribution, as in the two-phase split computing framework (TSCF) [ko2024twophase], or focusing on individual weights. Applied during or after training, dynamic and static pruning balance accuracy, and computational cost. For instance, Hybrid SD [yan2024hybrid] employs structural pruning of the Stable Diffusion Model’s U-Net to maintain image quality on edge devices, while the adaptive pruning-split federated learning (PSFed) method [zheng2024semantic] optimizes semantic coder updates in a SemCom-assisted SEC framework, reducing training delay and energy consumption.